OpenStudy (anonymous):

Factoring Trinomials: Would This Be The Correct Answer To 3x^2-7x+2? Answer: (x-2y) (3x-1)

OpenStudy (whpalmer4):

Here's how we find out: multiply! \[(x-2y)(3x-1) = x(3x-1) - 2y(3x-1) = 3x^2 -x -6xy + 2y \] Does that look like \(3x^2-7x+2\) to you? Where'd that \(y\) come from, anyhow?

OpenStudy (anonymous):

You Don't Have To Be Rude, I'm Not The One Who Answered This Question In The First Place.